OpenStudy (anonymous):

which of the following are square roots of the number below 100 a. -10 b.-(100)^1/2 c.100^1/2 d. 25 help please an thank u ?

OpenStudy (anonymous):

b and c are another way of writing square root of 100 or square root of -100. That doesn't mean it's an answer though. I would chose a for the answer.

OpenStudy (anonymous):

\[\LARGE \sqrt{x}=\left(x\right)^{\frac{1}{2}}\]

OpenStudy (anonymous):

-|x| is not a square root of x^2.

OpenStudy (anonymous):

Was that directed at me? -10 x -10 = 100

OpenStudy (anonymous):

\[\LARGE -(10)^2\neq 100\]

OpenStudy (anonymous):

@cococupcoffee you mentioned option ... and that is not true !

OpenStudy (anonymous):

option b

OpenStudy (anonymous):

sorry, what? i am confused lol

OpenStudy (anonymous):

A common mistake is to think that the square root is the inverse of squaring. This is only true for positive values of x.

OpenStudy (anonymous):



OpenStudy (anonymous):

No, sqrt(100) =10

OpenStudy (anonymous):

you said cococupcoffee 1 Good Answer b and c are another way of writing square root of 100 or square root of -100. That doesn't mean it's an answer though. I would chose a for the answer. and option b says b.-(100)^1/2 \[\LARGE -(100)^{1/2}=-\sqrt{100}=-10\] and we know that... \[\LARGE \sqrt{100}\neq -10\]

OpenStudy (anonymous):

really? i was taught something else in school then!

OpenStudy (anonymous):

\[\LARGE x^2=4 \Longrightarrow x=\pm\sqrt4\] but... \[\LARGE \sqrt9 \neq \pm 3\]

OpenStudy (anonymous):

It's confusing, I know.

OpenStudy (anonymous):

So, kreshnik, i kinda get what you're saying but, doesn't -3^2 = 9?

OpenStudy (anonymous):

@cococupcoffee only when we have X,Y, or what ever you can say plus minus because... \[\LARGE x=\pm \sqrt4 \] \[\LARGE x=2 \quad \quad x=-2\] because.. \[\LARGE x^2=4 \longrightarrow (-2)^2=4\]

OpenStudy (anonymous):

only when you have bracket !! like this ! : \[\LARGE (-3)^2=9 \quad \quad but...\quad \quad -3^2\neq9\]

OpenStudy (anonymous):

Oh... i am going to sound dumb... but since it's just -3 you're squaring, do you really need the bracket?

OpenStudy (anonymous):

@cococupcoffee It's not shame not to know! It's shame not wanting to know!! (Sokrat) Yes we need them because any number it is squared in POSITIVE number is equal + if it is odd then is MINUS ... (I mean when we have a minus number on square )..Here are some examples... \[\LARGE -1^2=-1 \quad ,\quad (-1)^2=1 \quad \text{Because...}\] \[\LARGE -1^2=-1\cdot 1=-1^2=-1\] and \[\LARGE (-1)^2=-1\cdot(-1)=+1\] \[\LARGE (-1)^{1999}=-1 \quad ,\quad (-1)^{2000}=1\]

OpenStudy (anonymous):

Do you understand it ?

OpenStudy (anonymous):

I get the odd and even part lol. I even get the part where you are saying that having a bracket means -n x -n, but wihtout bracket, it's -n x n. i am just surprised that i never learned that in school. thanks for the explanation :)

OpenStudy (anonymous):

here is why minus... \[\LARGE -1^2=-1\cdot +1=-1 \quad \quad \because...\] \[\LARGE (-)\cdot(-)=+ \quad ,\quad (+)\cdot(+)=+ \quad\] \[\LARGE (-)(+)=- \quad ,\quad (+)(-)=-\]

OpenStudy (anonymous):

lol i knew that :) ill ask my math teacher about the whole brackets thing to understand why they had to make it so complicated lol

OpenStudy (anonymous):

... ask your teacher, I can't explain them better, but you'll see that they are very easy ;)...

OpenStudy (anonymous):

I have to go.. see you later ;)

OpenStudy (anonymous):

and once again, thanks for pointing that out! n what do u think is the answer for this question

OpenStudy (anonymous):

alritie bye! :)

OpenStudy (anonymous):

\[\LARGE \sqrt x=(x)^{\frac12} \longrightarrow \sqrt{100}=\left(100\right)^{\frac12}\] C 100% bye !